Let \(p_1\) be the proportion of successes in the first population and let \(p_2\) be the proportion of successes in the second population. Suppose that you are testing the hypotheses:
\(H_0:p_1-p_2=0 \\ H_a:p_1-p_2 \neq 0\)
Futhermore suppose that \(z^*=1.73\), find and input the p-value for this test

Answer by stanbon(75887)

Let \(p1\) be the proportion of successes in the first population and let \(p2\) be the proportion of successes in the second population. Suppose that you are testing the hypotheses:
H_0:p1-p2=0
H_a:p1-p2 # 0
Futhermore suppose that \(z^*=1.73\), find and input the p-value for this test
----
p-value = 2*p(z > 1.73) = 2*normalcdf(1.73,100) = 0.0836
---------------
Cheers,
Stan H.
---------